### v1

# Momentum-space Metric Tensor from Nonadiabatic Evolution of Bloch Electrons

Abstract
We reveal a fundamental geometric structure of momentum space arising from the nonadiabatic evolution of Bloch electrons. By extending semiclassical wave packet theory to incorporate nonadiabatic effects, we introduce a momentum-space metric tensor—the nonadiabatic metric. This metric gives rise to two velocity corrections, dubbed geometric and geodesic velocities, providing a unified and intuitive framework for understanding nonlinear and nonadiabatic transport phenomena beyond Berry phase effects. Furthermore, we show that the nonadiabatic metric endows momentum space with a curved geometry, recasting wave packet dynamics as forced geodesic motion. In this picture, the metric defines distances, the Berry connection acts as a gauge potential, band dispersion serves as a scalar potential, and the toroidal topology of the Brillouin zone imposes periodic boundary conditions. When the nonadiabatic metric is constant, it reduces to an effective mass, allowing electrons to behave as massive particles in flat bands. In a flat Chern band with harmonic attractive interactions, the two-body wave functions mirror the Landau-level wave functions on a torus.
The adiabatic approximation, pioneered by Born and Oppenheimer, is a cornerstone of condensed matter physics. The emergence of Berry phase from adiabatic evolution [1] has profoundly shaped the study of topological phases of matter [2]. Adiabatic wave packet dynamics linked the Berry curvature to anomalous velocity in solids [3, 4, 5, 6], provided a unified geometric framework to understand topological phenomena [2]. This framework has revolutionized our understanding of how topological quantum materials respond to electric, magnetic, thermal, and lattice deformation stimuli [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Recent interest has surged in the Bloch-electron dynamics beyond Berry phase effects manifested in, for example, nonlinear responses [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. These phenomena are thought to stem from the quantum metric [23, 24, 25]—the real counterpart of the Berry curvature in the quantum geometric tensor [37]. The quantum metric is connected to rich physical effects, such as massive particle dynamics in flat bands [38, 39, 40, 41, 42, 43, 44, 45], exciton spectra [46, 47], fractional Chern insulators [48, 49, 50, 51], Wannier function spread [52], orbital magnetic susceptibility [53, 54], among others [23, 24, 25]. Despite its broad relevance, an intuitive framework for understanding those phenomena is still lacking.
In this Letter, we deliver a semiclassical perspective of the above phenomena by studying the nonadiabatic effects—the leading-order corrections to the Berry phase effects. While nonadiabatic effects have been explored in quantum systems driven by external classical parameters [55, 56, 57], their significance in the context of Bloch electrons remains elusive. Unlike the quantum-classical hybrid systems, Bloch electrons form a genuine quantum system, free from external classical degrees of freedom. Moreover, Bloch wavefunctions are parameterized by crystal momentum within the Brillouin zone—a torus characterized by nontrivial topology. This topological structure not only gives rise to the Berry phase in one dimension, known as the Zak phase [58], but also imposes periodic boundary conditions on electron dynamics within the parameter space, as we will demonstrate later.
Here, we extend semiclassical wave packet theory to incorporate nonadiabatic effects by including interband contributions (see Fig. 1). We begin by deriving the effective Lagrangian for the wave packet, revealing that nonadiabatic corrections emerge as a metric tensor in momentum space, which we term the nonadiabatic metric. We then discuss the relationship between this metric and the quantum metric and elucidate its profound implications for wave packet velocity, which provides a unified understanding of nonlinear and nonadiabatic transport. Next, we demonstrate that the nonadiabatic metric endows the Brillouin zone with the structure of a curved manifold, causing electron dynamics to resemble forced geodesic motion on this curved space. We further requantize the electron’s dynamics in momentum space, where the toroidal geometry of the Brillouin zone imposes periodic boundary conditions. When applied to flat Chern bands, this approach reveals the reemergence of massive particle dynamics and establishes a direct connection between bound-state wave functions and quantum Hall effects on a torus.
Figure 1: Illustration of nonadiabatic time evolution of a wave packet in momentum space, where a gradual momentum shift induces interband mixings and nonadiabatic corrections.
Our results indicate that, in the dynamics of the Bloch electrons, the nonadiabatic metric is more fundamental than the quantum metric: it emerges directly from the Schrödinger equation as the leading correction beyond adiabatic evolution. It governs the dynamical behavior of Bloch electrons, providing a unified description of various dynamical phenomena. In contrast, the quantum geometric tensor is defined in the adiabatic limit, where it describes the geometric properties of the static Hilbert space [37].
Nonadiabatic Wave Packet Theory— We use the following wave packet ansatz [59]:
[…]
$|\tilde{u}_{\bm{q}}\rangle=\sum_{n}c_{n}(\boldsymbol{q},t)|u_{n}(\boldsymbol{q})\rangle$ (2)
with $|u_{n}(\boldsymbol{q})\rangle$ the periodic part of the Bloch state for band $n$, $c_{n}$ the coefficient, and $\sum_{n}|c_{n}(\boldsymbol{q},t)|^{2}=1$. The wave packet center is at $\bm{x}_{c}=\langle\Psi|\hat{\bm{x}}|\Psi\rangle$.
Here we consider an isolated band labeled by index 0. We focus on the abelian case with $|c_{0}|\approx 1$ and $c_{n}\sim 0$ for $n\neq 0$. In contrast to the adiabatic wave packet dynamics developed by Sundaram and Niu [5], which constrains the wave packet on a single energy band, here we allow for interband contributions with $c_{n}\neq 0$ (see Fig. 1). These interband components are expressed as an asymptotic series in the frequency of external perturbations—a notable distinction from the higher-order theory in Ref. [27], where the wave packet is expanded as an asymptotic series in the strength of external fields.
The dynamics of the wave packet is obtained by minimizing the Dirac-Frenkel action [60, 61]
[…]
where $\bm{x}_{c}$, $\bm{q}_{c}$, $c_{n}$, and $c_{n}^{*}$ are dynamical variables. By integrating out the dynamics of $c_{n}$ [59], we obtain our central result: the nonadiabatic effective Lagrangian (with $\hbar=1$)
$\mathcal{L}_{\rm eff}$ $=\mathcal{L}_{\rm eff,0}+\frac{1}{2}G_{ij}\dot{q}_{c}^{i}\dot{q}_{c}^{j}$ (4)
$\mathcal{L}_{\rm eff,0}$ $=-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{0}(\bm{q}_{c})-V(\bm{x}_{c})+\dot{q}_{c}^{i}A_{i}$
where $\mathcal{L}_{\rm eff,0}$ is the adiabatic effective Lagrangian with ${A}_{i}=\langle u_{0}|i\partial_{{q}^{i}}|u_{0}\rangle|_{\bm{q}_{c}}$ the Berry connection, $E_{n}$ the eigenenergy of the $n$-th band, and $V(\bm{x}_{c})$ the slowly varying external potential. The Einstein summation convention is implied throughout the manuscript unless otherwise noted. From now on, we drop the subscript $c$, using $(\bm{x},\bm{q})$ for $(\bm{x}_{c},\bm{q}_{c})$.
Our key contribution is the identification of a new term in the effective Lagrangian, characterized by the coefficients $G_{ij}$. $G_{ij}$ acts as a metric tensor in momentum space, which is termed the nonadiabatic metric, as this metric originates from nonadiabatic corrections to wave packet dynamics: $c_{n}\propto\dot{q}^{i}$ for $n\neq 0$, which reflect a finite occupation of other instantaneous eigenstates. The matrix element reads
$G_{ij}=$ $2\text{Re}\sum_{m\neq 0}\frac{A_{i,[0m]}A_{j,[m0]}}{\omega_{[m0]}}$ (5)
where $A_{i,[mn]}=\langle u_{m}|i\partial_{q^{i}}|u_{n}\rangle$, $\omega_{[mn]}=E_{m}-E_{n}$. The metric is a gauge-invariant symmetric tensor [62]. The nonadiabatic metric is related to but different from the quantum metric $g_{ij}=\text{Re}\mathcal{Q}_{ij}$ that is the real part of the quantum geometric tensor (or the Fubini-Study metric) $\mathcal{Q}_{ij}=\sum_{m\neq n}{A_{i,[nm]}A_{j,[mn]}}$. In a two-band model with an energy gap $\Delta(\bm{q})$, $G_{ij}$ of the lower band is proportional to $g_{ij}$: $G_{ij}=\frac{2}{\Delta}g_{ij}$.
Just as adiabatic theory unifies Berry phase effects through the term $\dot{q}^{i}A_{i}$ [2], the nonadiabatic theory introduces the new term, $\frac{1}{2}G_{ij}\dot{q}^{i}\dot{q}^{j}$, providing a comprehensive framework for capturing a wide array of phenomena that extend beyond the Berry phase effects. As we show below, it is the nonadiabatic metric, rather than the quantum metric, that fundamentally governs a broad class of nonlinear and nonadiabatic transport phenomena.
At a deeper level, the nonadiabatic metric uncovers the intrinsically curved geometric structure of the Brillouin zone ($\bm{q}$-space) and highlights its topological character as a torus, as detailed later. Notably, this $\bm{q}$-space representation enables the reemergence of massive particle dynamics in flat bands with corrections from interband mixing, connecting the bound states of particles in flat Chern bands to quantum Hall effects on a torus.
Nonlinear and Nonadiabatic Transport— The nonadiabatic metric shows a profound impact on the electronic transport phenomena by modifying the wave packet velocity $\dot{x}_{i}$ in the following set of equations of motion [59]:
$\dot{q}^{i}$ $=-\frac{\partial V}{\partial x_{i}}$ (6)
$\dot{x}_{i}$ $=\frac{\partial E}{\partial q^{i}}-F_{ij}\dot{q}^{j}+\Gamma_{i,jk}\dot{q}^{j}\dot{q}^{k}+G_{ij}\ddot{q}^{j}$ (7)
where the velocity $\dot{x}_{i}$ has four different contributions. The first term is the group velocity from band dispersion that was derived back to the 1930s [63, 5]. The second term is the anomalous velocity from the Berry curvature $F_{ij}=-2\text{Im}\mathcal{Q}_{ij}$ arising from the adiabatic evolution of wave packets advanced by Chang, Sundaram, and Niu [7, 8, 5]. The nonadiabatic metric leads to two velocity corrections, termed the geodesic velocity and geometric velocity separately, as detailed below.
The kinetic energy of the generalized coordinate $\bm{q}$ leads to a contribution termed the geodesic velocity, as the coefficient $\Gamma_{i,jk}$ is closely related to $\Gamma^{i}_{jk}$ that defines the geodesic motion as shown later in Eq. (10). $\Gamma_{i,jk}$ is the Christoffel symbol of the first kind:
$\Gamma_{i,jk}$ $=\frac{1}{2}\left(\frac{\partial G_{ij}}{\partial q^{k}}+\frac{\partial G_{ik}}{\partial q^{j}}-\frac{\partial G_{jk}}{\partial q^{i}}\right).$ (8)
In the presence of a static electric field $\bm{\varepsilon}$, $\dot{\bm{q}}=\bm{\varepsilon}$ and $\ddot{\bm{q}}$ vanishes. The velocity thus reduces to
$\dot{x}_{i}=\frac{\partial E}{\partial q^{i}}-F_{ij}\varepsilon^{j}+\Gamma_{i,jk}\varepsilon^{j}\varepsilon^{k},$ (9)
which has been widely adopted in studying nonlinear responses [27, 64, 65, 23, 24, 25]. The velocity in Eq. (9) was originally derived by studying adiabatic evolution [27, 64], where the wave-packet wave function is expanded as an asymptotic series in the field strength. In contrast, our derivation demonstrates that the same correction can be systematically obtained by expanding the wave function as an asymptotic series in the frequency of the external fields. This perspective naturally generalizes the framework, enabling the analysis of electronic responses to a broader range of perturbations beyond static electric and magnetic fields.
The second nonadiabatic velocity is new, termed the geometric velocity, as the coefficient is the nonadiabatic metric. This term is proportional to $\ddot{\bm{q}}$, the acceleration of the generalized coordinate $\bm{q}$. While $\ddot{\bm{q}}$ vanishes under static fields or is negligible in the adiabatic limit, it becomes significant in nonadiabatic responses to time-dependent, oscillating forces. For instance, in the presence of a finite-frequency electric field $\varepsilon^{i}e^{i\omega t}$, the geometric velocity acquires a nonzero value: $\ddot{q}^{i}=i\omega\varepsilon^{i}e^{i\omega t}$. At linear order in $\varepsilon^{i}$, this term gives rise to a dissipationless polarization current, which agrees with the linear response theory based on the Kubo formula [65, 66]. The full implication of the geometric velocity remains an open area for further investigation. It will be manifested in nonlinear and nonadiabatic responses to finite-frequency driving, such as below-gap light [67, 68, 69, 70] and coherent collective excitations like magnons and phonons, which are phenomena of broad relevance [71, 72, 73, 74, 75, 76].
This unified framework emphasizes the fundamental role of the nonadiabatic metric $G_{ij}$. While earlier work associated $G_{ij}$ with Berry connection polarizability via expansions in field strength [27, 64, 65], our results establish its inherently nonadiabatic character, emerging from expansions in field frequency. This perspective not only provides a unified and complete understanding of nonlinear and nonadiabatic responses but also naturally generalizes to responses to non-electromagnetic drivings, such as magnonic, phononic, or thermal fields—where the nonadiabatic metric $G_{ij}$ likewise plays a pivotal role.
$\bm{q}$-Space Picture: Metric Tensor and Forced Geodesic Equation— Here we unveil the deeper geometric significance of $G_{ij}$ in momentum space, or $\bm{q}$-space, to highlight the role of $\bm{q}$ as a generalized coordinate. $G_{ij}(\bm{q})$ in the Lagrangian defines a metric tensor in the tangent vector space of the generalized velocity $\dot{\bm{q}}$ defining a line element $ds^{2}=G_{ij}dq^{i}dq^{j}$. Different from the quantum metric that defines the distance between wavefunctions in the momentum space [37], here the distance is defined for the generalized coordinate $\bm{q}$ describing the wave packet center. A $\bm{q}$-dependent $G_{ij}$ signifies a curved manifold. Conversely, in cases where $G_{ij}$ is constant across the momentum space, it corresponds to an effective mass in a flat $\bm{q}$-space, simplifying the dynamics to a Euclidean framework.
Expressing the equations of motion in the $\bm{q}$-space clarifies the geometric interpretation. When the inverse of the metric tensor exists, the equation of motion reads
$\ddot{q}^{i}+\Gamma^{i}_{jk}\dot{q}^{j}\dot{q}^{k}=$ $G^{ij}\left(-\frac{\partial E}{\partial q^{j}}+F_{jk}\dot{q}^{k}+\dot{x}_{j}\right)$ (10)
where $\Gamma^{l}_{jk}$ is the Christoffel symbol of the second kind
$\Gamma^{i}_{jk}=\frac{1}{2}G^{il}\left(\frac{\partial G_{lj}}{\partial q^{k}}+\frac{\partial G_{kl}}{\partial q^{j}}-\frac{\partial G_{jk}}{\partial q^{l}}\right),$ (11)
$G^{ji}G_{ik}=\delta^{j}_{k}$ with $\delta^{j}_{k}$ the Kronecker delta. When the right-hand side of Eq. (10) vanishes, the equation reduces to the geodesic equation for a free particle navigating a curved manifold with metric tensor $G_{ij}$. A nonzero right-hand side indicates a particle subject to forces from three sources (see Fig. 2): a scalar potential from the band dispersion $E$, a gauge field from the Berry curvature $F_{ij}$, and an external force from the wave packet velocity $\dot{x}_{j}$. This equation enables the study of gravity in condensed matter setups [77]. It is noted that Eq. (10) differs from the recent study on momentum space gravity [78, 79, 80] where $G^{ij}$ and $\ddot{\bm{q}}$ are absent and $\Gamma_{i,jk}$ instead of $\Gamma^{i}_{jk}$ is used.
Figure 2: Wave packet dynamics in $\bm{q}$-space, a torus with curved geometry. The distance $ds^{2}$ is defined by the nonadiabatic metric $G_{ij}$. The topological charge stands for the Dirac monopole that generates nonzero gauge fields corresponding to the Berry curvature. $C$ indicates the Chern number. Color encodes the scalar potential from the band dispersion.
When $x_{i}$ can be invertibly represented as a function of $\dot{q}^{j}$ by using $\dot{q}^{i}=-\frac{\partial V}{\partial x_{i}}$, we have
$x_{i}=x_{i}(\dot{\bm{q}}),~{}~{}~{}\dot{x}_{i}=\frac{\partial x_{i}}{\partial\dot{q}^{j}}\ddot{q}^{j}=-\Xi_{ij}\ddot{q}^{j},$ (12)
the equations of motion find their $\bm{q}$-space representation:
$\ddot{q}^{i}+\bar{\Gamma}^{i}_{jk}\dot{q}^{j}\dot{q}^{k}=$ $\bar{G}^{ij}\left(-\frac{\partial E}{\partial q^{j}}+F_{jk}\dot{q}^{k}\right),$ (13)
where $\bar{G}_{ij}=G_{ij}+\Xi_{ij}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol. With information of $\dot{\bm{q}}$, $x_{i}$ can be solved using algebraic equations. The metric tensor becomes dynamical with $\dot{\bar{G}}_{ij}=\frac{\partial G_{ij}}{\partial q_{k}}\dot{q}_{k}+\frac{\partial\Xi_{ij}}{\partial\dot{q}_{k}}\ddot{q}_{k}$.
A concrete example is the harmonic potential
$V(\bm{x})=\frac{\kappa}{2}\bm{x}^{2},$ (14)
which leads to $x^{i}=-\frac{1}{\kappa}\dot{q}^{i}$ and $\dot{x}^{i}=-\frac{1}{\kappa}\ddot{q}^{i}$. Here $x^{i}=\delta^{ij}x_{j}$ and $\delta^{ij}$ the Kronecker delta. The equation of $\bm{q}$ is that in Eq. (13) with $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$. This equation follows from the Lagrangian [59]
$L=\frac{1}{2}\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}-A_{i}\dot{q}^{i}-E(\bm{q}),$ (15)
where $E(\bm{q})$ is a scalar potential and $A_{j}$ is a vector potential. Unlike a relativistic particle in curved spacetime, where $\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}=\text{constant}$ due to the constant seed of light, no such constraint applies here.
This Lagrangian has a well-defined canonical structure enabling us to perform canonical quantization. The canonical momentum is
$p_{i}=\frac{\partial L}{\partial\dot{q}^{i}}=\bar{G}_{ij}\dot{q}^{j}-A_{i}.$ (16)
In specific, with a constant nonadiabatic metric $G_{ij}=G_{0}\mathcal{I}_{ij}$ with $\mathcal{I}$ the identity matrix, an effective mass emerges and the Hamiltonian simplifies to
$H=\frac{1}{2M}\delta^{ij}(p_{i}+A_{i})(p_{j}+A_{j})+E(\bm{q})$ (17)
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. The covariant Laplacian will need to be used when $G_{ij}$ is not a constant.
By promoting $(p_{i},q^{j})$ to operators with the commutator $[\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian. The re-quantization of semiclassical wave packet dynamics is fruitful in describing various quantum phenomena such as Landau levels of Bloch electrons [6] and Berry curvature effects on exciton spectrum [47]. We note that here ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 47, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This contrasts with previous approaches [6, 47], which performed re-quantization in real-space coordinates without imposing such conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus, as discussed below.
In the presence of a time-dependent uniform electric field, the harmonic potential acquires a time-dependent center, leading to a time-dependent vector potential in the effective Hamiltonian that encodes the influence of the electric field.
This $\bm{q}$-space framework provides an alternative lens for studying Bloch electron dynamics under static or dynamic confining potentials, a class of problems that are particularly interesting in flat-band systems, where finite mass and geodesic motion arise from the interplay between the potential and the metric $G_{ij}$.
Single Particle Dynamics in a Flat Band— The scalar potential $E(\bm{q})$ is a constant in flat bands, and the electron dynamics are determined by the Berry curvature and nonadiabatic metric. Here we study a two-dimensional flat band. We focus on the simplest case when the Berry curvature $F_{ij}$ and nonadiabatic metric $G_{ij}$ are constants, and leave the non-uniform case for future study. This simplifies the dynamics from a curved manifold to a Euclidean torus. Consider a rectangular Brillouin zone with dimensions $L_{q_{x},q_{y}}=2\pi/L_{x,y}$ where $L_{q_{x},q_{y}}$ defines the ${\bm{q}}$-space size while $L_{x,y}$ are the real-space unit cell dimensions.
When the particle in this flat band is confined in a harmonic potential shown in Eq. (14), the Hamiltonian in $\bm{q}$-space follows Eq. (17)
$\hat{H}=\frac{1}{2M}(\hat{\bm{p}}+\hat{\bm{A}})^{2},$ (18)
where $M=G_{0}+\frac{1}{\kappa}$ for a constant metric $G_{ij}=G_{0}\mathcal{I}_{ij}$, and $\hat{\bm{A}}=\frac{1}{2}F\hat{z}\times\hat{\bm{q}}$, corresponding to $F_{12}=-F_{21}=-F$. This describes a Schrödinger equation with periodic boundary conditions in $\bm{q}$-space.
The Berry curvature’s integral over the Brillouin zone corresponds to the Chern number $C$, i.e., $FL_{q_{x}}L_{q_{y}}=2\pi C$. This indicates a quantized effective magnetic flux generated by $C$ Dirac monopoles within the torus. This ensures the self-consistency of the equation, i.e., the existence of solutions satisfying the boundary condition.
With nonzero $C$, the eigenvalue problem of Eq. (18) becomes a Landau level problem on a torus [81, 82]. The eigenfunctions are Landau level wavefunctions on a torus. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F\kappa}{1+\kappa G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass, whereas the nonadiabatic metric contributes to a correction from interband mixing. The interband mixing effect is enhanced by increasing the potential strength that increases $\dot{\bm{q}}$ and thus the interband contributions.
Alternative to TKNN’s approach, which employs a periodic potential to lift Landau level degeneracy [83], a harmonic potential can also lift the degeneracy of a flat band. This maps the flat-band particle within a confining potential in real space into a Landau level problem on a torus in $\bm{q}$-space, allowing for analytical solutions of both eigenfunctions and eigenenergies. The nonadiabatic metric induces corrections from interband mixing. The metric contribution is significant when $|\kappa G_{0}|\gg 1$.
Figure 3: Bound states of flat-band particles in real space mirror the Landau-level wave functions on a torus in $\bm{q}$-space.
Two-body Bound States in Flat Bands— The effect of $G_{0}$ can be enhanced in a two-body problem [47] where $\kappa$ represents the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the wave packets are well separated, and they interact through a softened Coulomb interaction in real space. The two-particle Lagrangian is
$L_{\rm T}=L({\bm{x}}_{1})+L({\bm{x}}_{2})+\frac{1}{\epsilon\sqrt{(\bm{x}_{1}-\bm{x}_{2})^{2}+x_{0}^{2}}}$ (19)
where $\bm{x}_{1,2}$ are wave packet centers, $L(\bm{x}_{1,2})$ is the single-particle Lagrangian with the same constant Berry curvature $F$ and nonadiabatic metric $G_{0}$, $\epsilon$ is the effective dielectric constant, the charge is unity, and $x_{0}$ is the cutoff length. Focusing on bound states with $|\bm{x}_{1}-\bm{x}_{2}|\ll x_{0}$, we approximate the Lagrangian as
$L_{\rm T}\simeq L({\bm{x}}_{1})+L({\bm{x}}_{2})+\left[V_{0}-\frac{1}{2}\kappa_{\rm I}({\bm{x}}_{1}-{\bm{x}}_{2})^{2}\right]$ (20)
where $V_{0}=\frac{1}{\epsilon x_{0}}$ and $\kappa_{\rm I}=\frac{V_{0}}{x_{0}^{2}}$.
We focus on the relative dynamics. Applying canonical quantization, the effective Hamiltonian for the relative coordinate is
$\hat{H}_{\rm-}=\frac{1}{2M^{\prime}}(\hat{\bm{p}}_{-}+\hat{\bm{A}}(\hat{\bm{q}}_{-}))^{2}-V_{0}$ (21)
where $\hat{\bm{q}}_{-}=(\hat{\bm{q}}_{1}-\hat{\bm{q}}_{2})/\sqrt{2}$, $[\hat{p}_{-,i},\hat{q}_{-}^{j}]=-i\delta_{i}^{j}$, and $M^{\prime}=G_{0}+\frac{1}{2\kappa_{\rm I}}$ is the reduced effective mass. This Hamiltonian again resembles a Landau level problem on a torus (see Fig. 3) with eigenenergies $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M^{\prime}}-V_{0}$.
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. If they are indistinguishable, the two-body wavefunctions need to be antisymmetric for fermions or symmetric for bosons. This leads to another boundary condition in addition to the periodic boundary conditions. Therefore, while the degeneracy of eigen-energy levels obtained by solving Eq. (21) is the Chern number $|C|$ ($C\neq 0$), only those states satisfying both boundary conditions are allowed solutions.
A natural realization of the two-body problem is the excitons formed by electrons and holes in flat bands [84, 47, 85]. The potential relationship with Cooper pairs is interesting and warrants further investigation [86, 87, 88, 89].
Summary— We introduce a geometric framework that extends the study of Bloch electron dynamics beyond adiabatic Berry-phase effects. We show that the leading-order corrections to the adiabatic dynamics, the nonadiabatic effects, are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor in semiclassical wave packet dynamics. This metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature, enabling further generalization. Fundamentally, the nonadiabatic metric endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space. The momentum space picture provides an intuitive understanding of bound-state problems of flat-band particles, where massive dynamics is retrieved and the nonadiabatic metric captures interband mixing effects. Notably, the two-body bound-state wave functions in flat Chern bands resemble Landau level states on a torus.
Acknowledgements.
Acknowledgments— Y.R. is indebted to Qian Niu, Yang Gao, Di Xiao, Wenqin Chen and Jie Wang for useful discussions.