### v2

# Nonadiabatic Origin of Quantum-Metric Effects via Momentum-Space Metric Tensor

Abstract
We reveal a fundamental geometric structure of momentum space arising from the nonadiabatic evolution of Bloch electrons. By extending semiclassical wave packet theory to incorporate nonadiabatic effects, we introduce a momentum-space metric tensor—the nonadiabatic metric. This metric gives rise to two velocity corrections, dubbed geometric and geodesic velocities, providing a unified and intuitive framework for understanding nonlinear and nonadiabatic transport phenomena beyond Berry phase effects. When the nonadiabatic metric is constant, it reduces to an effective mass, modifying flat-band electron dynamics in confining potentials. In a flat Chern band with harmonic attractive interactions, the two-body wave functions mirror the Landau-level wave functions on a torus. Furthermore, we show that the nonadiabatic metric endows momentum space with a curved geometry, recasting wave packet dynamics as forced geodesic motion.
I Introduction
The adiabatic approximation, pioneered by Born and Oppenheimer, is a cornerstone of condensed matter physics. The emergence of Berry phase from adiabatic evolution [1] has profoundly shaped the study of topological phases of matter [2]. Recent interest has surged in the effects of quantum geometry in Bloch-electron dynamics beyond the Berry-phase effects, which are thought to stem from the quantum metric [3, 4, 5, 6]—the real counterpart of the Berry curvature in the quantum geometric tensor [3]. The quantum metric is connected to rich and diverse physical effects, such as nonlinear responses [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], massive particle dynamics in flat bands [21, 22, 23, 24, 25, 26, 27, 28], exciton spectra [29, 30], fractional Chern insulators [31, 32, 33, 34], Wannier function spread [35], orbital magnetic susceptibility [36, 37], among others [4, 5, 6]. Pinpointing the key physical processes in which quantum geometric effects play a significant role is essential for developing a unified understanding of their diverse manifestations.
In this work, we reveal the fundamental connection between the effects of quantum geometry and the nonadiabatic electronic dynamics. We study the nonadiabatic Bloch electron dynamics by using the semiclassical wave packet theory. The wave packet theory has connected the adiabatic Berry curvature to anomalous velocity of Bloch electrons [38, 39, 40, 41], providing a unified framework to understand Berry-phase effects in crystalline materials [2] and revolutionizing our understanding of how topological quantum materials respond to electric, magnetic, thermal, and lattice deformation stimuli [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57].
We extend the adiabatic wave packet theory to the nonadiabatic regime by including interband contributions (see Fig. 1). We begin by deriving the effective Lagrangian for the wave packet, revealing that nonadiabatic corrections emerge as a metric tensor in momentum space, which we term the nonadiabatic metric, in addition to other corrections. The relationship between this metric and the quantum metric is discussed. In the following, we elucidate on the profound implications of the nonadiabatic metric. We first discuss its impact on the wave packet velocity, which provides a unified understanding of nonlinear and nonadiabatic transport. Then we explore its effect on the quantum dynamics of Bloch electrons in flat-band systems, highlighting the interband mixing-induced correction on the bound-state spectrum.
Our results indicate that, in the dynamics of the Bloch electrons, the nonadiabatic metric is more fundamental than the quantum metric: it emerges directly from the Schrödinger equation as the leading correction beyond adiabatic evolution. It governs the dynamical behavior of Bloch electrons, providing a unified description of various dynamical phenomena. In contrast, the quantum geometric tensor is defined to describe the geometric properties of static wave functions on a parameter space, which are naturally tied to adiabatic processes [3].
Figure 1: Illustration of nonadiabatic time evolution of a wave packet in momentum space, where a gradual momentum shift induces interband mixings and nonadiabatic corrections.
II Results
II.1 Formalism: Nonadiabatic Wave Packet Theory
We study the dynamics of a wave packet centered at $\bm{x}_{c}$ under slowly varying scalar potential $V(\bm{x})$. Its spatial spread is much smaller than the perturbation scale, which enables us to consider an approximated Hamiltonian local to the wave packet. The Hamiltonian can be approximated by linearizing perturbations around the wave packet center:
$\mathcal{H}\approx$ $\hat{H}_{c}+\hat{H}_{1},$ (1)
where $\hat{H}_{c}=\hat{H}(\hat{\bm{x}},\hat{\bm{p}})+V(\bm{x}_{c})$ represents the local Hamiltonian with $V(\bm{x}_{c})$ a constant and $\hat{H}$ the Hamiltonian operator of the unperturbed crystal. The corresponding Bloch bands and Bloch wave functions satisfying:
$\hat{H}|\psi_{n,\bm{q}}\rangle=E_{n}(\bm{q})|\psi_{n,\bm{q}}\rangle,$ (2)
$\hat{H}_{1}=(\hat{\bm{x}}-\bm{x}_{c})\cdot\nabla_{\bm{x}}V(\bm{x}_{c})$ is the gradient correction. Here $\nabla_{\bm{x}}V(\bm{x}_{c})$ are constants and do not induce inter-band coupling [40].
We use the following wave packet ansatz [58]:
$|\Psi(t)\rangle=\int d\bm{q}\,a(\boldsymbol{q},t)|\tilde{\psi}_{\boldsymbol{q}}\rangle,$ (3)
where $a(\boldsymbol{q},t)=|a|e^{-i\gamma}$ is a narrow distribution in momentum space centered at $\boldsymbol{q}_{c}=\int\bm{q}|a|^{2}d\bm{q}$, $|\tilde{\psi}_{\boldsymbol{q}}\rangle=e^{i\boldsymbol{q}\cdot\hat{\boldsymbol{x}}}|\tilde{u}_{\bm{q}}\rangle$ is a superposition of Bloch waves with quasimomentum $\bm{q}$, and
$|\tilde{u}_{\bm{q}}\rangle=\sum_{n}c_{n}(\boldsymbol{q},t)|u_{n}(\boldsymbol{q})\rangle$ (4)
with $|u_{n}(\boldsymbol{q})\rangle$ the periodic part of the Bloch wave function $|\psi_{n,\bm{q}}\rangle$, $c_{n}$ the coefficient, and $\sum_{n}|c_{n}(\boldsymbol{q},t)|^{2}=1$. The wave packet center is at $\bm{x}_{c}=\langle\Psi|\hat{\bm{x}}|\Psi\rangle$.
Here we consider an isolated band labeled by index 0. We focus on the abelian case with $|c_{0}|\approx 1$ and $c_{n}\sim 0$ for $n\neq 0$. In contrast to the adiabatic wave packet dynamics developed by Sundaram and Niu [40], which constrains the wave packet on a single energy band, here we allow for interband contributions with $c_{n}\neq 0$ (see Fig. 1). These interband components $c_{n}$ are subsequently expressed as an asymptotic series in the rate of parameter evolution such as $\dot{\bm{q}}_{c}$. This contrasts with the higher-order expansion in Ref. [8], where the wave packet is expanded as a series in the strength of external fields rather than in the rate of change of the underlying parameters.
The dynamics of the wave packet is obtained by minimizing the Dirac-Frenkel action (with $\hbar=1$[59, 60]
$\mathcal{S}$ $=\int dt\langle\Psi|i\frac{d}{dt}-\mathcal{H}|\Psi\rangle=\int dtL$ (5)
$L$ $=\sum_{n}c_{n}^{*}i\dot{c}_{n}-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}+\Lambda_{[00]}+\sum_{n\neq 0}-c_{n}^{*}c_{n}\left[\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]})\right]+(c_{0}^{*}c_{n}\Lambda_{[0n]}+c.c.)$
where $\dot{c}_{n}=dc_{n}/dt$, $\omega_{[mn]}=E_{m}-E_{n}$, $\Lambda_{[mn]}=\dot{\bm{q}}_{c}\cdot\bm{A}_{[mn]}$, and Berry connection is ${A}_{i,[mn]}=\langle u_{m}|i\partial_{{q}^{i}}|u_{n}\rangle|_{\bm{q}_{c}}$. Here $\bm{x}_{c}$, $\bm{q}_{c}$, $c_{n}$, and $c_{n}^{*}$ are dynamical variables. By writing down the equation of motion of $c_{n}$ and integrating out their dynamics [see Methods], we obtain our central result: the nonadiabatic effective Lagrangian
$\mathcal{L}_{\rm eff}$ $=-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}(\bm{q}_{c},\bm{x}_{c})+\dot{q}_{c}^{i}{A}_{i}+\frac{1}{2}G_{ij}\dot{q}_{c}^{i}\dot{q}_{c}^{j}$ (6)
where the Einstein summation convention is used for Cartesian coordinate indices throughout the manuscript, while summation over the band indices is written explicitly. Here $E_{c}=E_{0}(\bm{q}_{c})+V(\bm{x}_{c})$ with $E_{n}$ the eigenenergy of the $n$-th band and $V(\bm{x}_{c})$ the slowly varying external potential. ${A}_{i}$ is used as shorthand for ${A}_{i}^{00}$, capturing the Berry phase effects from adiabatic evolution of Bloch states. We emphasize that this work focuses on a slowly varying scalar potential, resulting in a gradient correction $\hat{H}_{1}$ that is proportional to a constant times $\hat{\bm{x}}-\bm{x}_{c}$. This specific form of $\hat{H}_{1}$ does not induce inter-band coupling. However, in the presence of a magnetic field or inhomogeneities, such as those arising from spin textures, the nonadiabatic theory yields additional corrections to the energy $E_{c}$ and Berry connections $A_{i}$ as discussed in our follow-up work [61]. From now on, we drop the subscript $c$, using $(\bm{x},\bm{q},E)$ for $(\bm{x}_{c},\bm{q}_{c},E_{c})$.
Our key contribution is the identification of a metric term beyond the Berry phase effects, characterized by the coefficients $G_{ij}$. $G_{ij}$ acts as a metric tensor in momentum space, which is termed the nonadiabatic metric, as this metric originates from nonadiabatic corrections to wave packet dynamics: $c_{n}\propto\dot{q}^{i}$ for $n\neq 0$, which reflect a finite occupation of other instantaneous eigenstates. The matrix element reads
$G_{ij}=$ $2\text{Re}\sum_{m\neq 0}\frac{A_{i,[0m]}A_{j,[m0]}}{\omega_{[m0]}}.$ (7)
The metric is a gauge-invariant symmetric tensor [62]. The nonadiabatic metric is related to but different from the quantum metric $g_{ij}=\text{Re}\mathcal{Q}_{ij}$ that is the real part of the quantum geometric tensor (or the Fubini-Study metric) $\mathcal{Q}_{ij}=\sum_{m\neq n}{A_{i,[nm]}A_{j,[mn]}}$. In a two-band model with an energy gap $\Delta(\bm{q})$, $G_{ij}$ of the lower band is proportional to $g_{ij}$: $G_{ij}=\frac{2}{\Delta}g_{ij}$, which is positive-semidefinite. For the upper band, the nonadiabatic metric is negative-semidefinite. The energy denominator makes $G_{ij}$ not necessarily positive-semidefinite and thus not necessarily a Riemannian metric.
Just as adiabatic wave packet theory unifies Berry phase effects through the term $\dot{q}^{i}A_{i}$ [2], the nonadiabatic corrections provide a comprehensive framework for capturing a wide array of phenomena that extend beyond the Berry phase effects. Below, we discuss the nonadiabatic effects in transport and quantum dynamics of Bloch electrons.
II.2 Nonlinear and Nonadiabatic Transport
Quantum-metric contributions have been identified in diverse electronic transport phenomena, ranging from nonlinear responses [4, 5, 6] to quantum capacitance [63]. Yet, in contrast to Berry-phase physics, where wave packet dynamics supply a unified and intuitive interpretive framework, no analogous picture exists for understanding the origin of these quantum-metric-related effects. Our nonadiabatic theory provides such a framework. Specifically, the nonadiabatic metric contributes directly to the wave-packet velocity, as expressed in the following equations [58]:
$\dot{q}^{i}$ $=-\frac{\partial E}{\partial x_{i}}$ (8)
$\dot{x}_{i}$ $=\frac{\partial E}{\partial q^{i}}-F_{ij}\dot{q}^{j}+\Gamma_{i,jk}\dot{q}^{j}\dot{q}^{k}+G_{ij}\ddot{q}^{j}$ (9)
where the velocity $\dot{x}_{i}$ has four different contributions. The first term is the group velocity from band dispersion that was derived back to the 1930s [64, 40]. The second term is the anomalous velocity from the Berry curvature $F_{ij}=\partial_{q^{i}}{A}_{j}-\partial_{q^{j}}{A}_{i}$ arising from the adiabatic evolution of wave packets advanced by Chang, Sundaram, and Niu [42, 43, 40]. In addition, the nonadiabatic metric leads to two new velocity corrections, termed the geodesic velocity and geometric velocity separately, as detailed below.
The nonadiabatic metric term behaves like a kinetic energy of the generalized coordinate $\bm{q}$ in a curved space. The first velocity correction from it is termed the geodesic velocity, as the coefficient $\Gamma_{i,jk}$ is closely related to $\Gamma^{i}_{jk}$ that defines the geodesic motion as shown later in equation (19). $\Gamma_{i,jk}$ is the Christoffel symbol of the first kind:
$\Gamma_{i,jk}$ $=\frac{1}{2}\left(\frac{\partial G_{ij}}{\partial q^{k}}+\frac{\partial G_{ik}}{\partial q^{j}}-\frac{\partial G_{jk}}{\partial q^{i}}\right).$ (10)
In the presence of a static electric field $\bm{\varepsilon}$, $\dot{\bm{q}}=\bm{\varepsilon}$ and $\ddot{\bm{q}}$ vanishes (i.e., we set the electron charge $e$ to unity). The velocity thus reduces to
$\dot{x}_{i}=\frac{\partial E}{\partial q^{i}}-F_{ij}\varepsilon^{j}+\Gamma_{i,jk}\varepsilon^{j}\varepsilon^{k},$ (11)
which has been widely adopted in studying nonlinear responses [8, 65, 63, 4, 5, 6]. By combining the velocity with the Boltzmann distribution function $f(\bm{q})=f_{0}(\bm{q}-\tau\dot{\bm{q}})$ under the relaxation time approximation, where $f_{0}$ is the Fermi-Dirac distribution function and $\tau$ is the relaxation time, one can calculate the current $j_{i}=\int\dot{x}_{i}(\bm{q})f(\bm{q})d\bm{q}$ and thereby the linear and nonlinear conductance. It is noted that the velocity in equation (11) was originally derived by studying adiabatic evolution [8, 65], using an asymptotic expansion in external field strengths. In contrast, our derivation uses an expansion in the time derivatives of intrinsic dynamical variables, which reproduces previous results in the static field limit.
The second nonadiabatic velocity is new, termed the geometric velocity, as the coefficient is the nonadiabatic metric $G_{ij}$. This term is proportional to $\ddot{\bm{q}}$, the acceleration of the generalized coordinate $\bm{q}$. While $\ddot{\bm{q}}$ vanishes under static fields or is negligible in the adiabatic limit, it becomes significant in nonadiabatic responses to time-dependent, oscillating forces. For instance, in the presence of a finite-frequency electric field $\varepsilon^{i}=\varepsilon_{0}^{i}e^{i\omega t}$ with $\varepsilon_{0}^{i}$ the amplitude and $\omega$ the frequency, we find that $\dot{q}^{i}={\varepsilon}^{i}$ and thus the geometric velocity acquires a nonzero value: $G_{ij}\ddot{q}^{i}=G_{ij}\dot{\varepsilon}^{i}$. At linear order in $\varepsilon^{i}_{0}$, this term gives rise to a dissipationless polarization current $j_{i}\sim\int G_{ij}f(\bm{q})d\bm{q}\dot{\varepsilon}^{j}$, which agrees with the linear response theory based on the Kubo formula (see Sec. III in the Supplemental Material [58][63, 66]. Specifically, in an insulator, the integral corresponds to the electric susceptibility $\chi_{ij}=\int G_{ij}d\bm{q}$ [63]. This is in contrast to the quantum metric, the integral of which in the Brillouin zone is related to the spread of the Wannier function [35]. The full implication of the geometric velocity remains an open area for further investigation. It will be manifested in nonlinear and nonadiabatic responses to finite-frequency driving, such as below-gap light [67, 68, 69, 70], second-harmonic generation [71], and coherent collective excitations like magnons and phonons, which are phenomena of broad relevance [72, 73, 74, 75, 76, 77].
This unified framework emphasizes the fundamental role of the nonadiabatic metric $G_{ij}$. While earlier work associated $G_{ij}$ with Berry connection polarizability via expansions in strengths of external fields [8, 65, 63], our results establish its inherently nonadiabatic nature, emerging from expansions in the changing rates of intrinsic dynamical parameters. This perspective not only offers a unified understanding of nonlinear and nonadiabatic responses, but also naturally extends to responses to non-electromagnetic fields, such as lattice dynamics and spin precessions, where the nonadiabatic metric $G_{ij}$ likewise plays a pivotal role, a feature not captured by previous theories.
II.3 Effective Mass Corrections for Bound States
The re-quantization of semiclassical wave packet dynamics is fruitful in describing various quantum phenomena such as Landau levels of Bloch electrons [41] and Berry curvature effects on exciton spectrum [30] and Wigner crystal phases [78]. While the importance of adiabatic Berry phase on those phenomena has attracted much attention in emerging material systems, such as transition-metal dichalcogenides, gapped graphenes, Landau levels, and moiré systems [30, 29, 79, 80, 78, 81, 82, 83], the effects of nonadiabatic corrections are yet to be explored.
We focus on the lowest Landau level as a concrete example. The simplicity enables analytical insights that can be generalized qualitatively to other flat-band systems. For the lowest Landau level, the Berry curvature $F_{ij}$ and nonadiabatic metric $G_{ij}$ are constants. Consider a rectangular Brillouin zone with dimensions $L_{q_{x},q_{y}}=2\pi/L_{x,y}$ where $L_{q_{x},q_{y}}$ defines the $\bm{q}$}-space size while $L_{x,y}$ are the dimensions of real-space unit cell with one magnetic flux quantum. The Berry curvature is uniform and its integral over the Brillouin zone corresponds to the Chern number $C$, i.e., $F=\frac{2\pi C}{L_{q_{x}}L_{q_{y}}}$. The nonadiabatic metric reads $G_{ij}=G_{0}\mathcal{I}_{ij}$ with $\mathcal{I}$ the identity matrix and $G_{0}=\frac{|F|}{\Delta}$ where $\Delta$ is the energy spacing between the lowest and the first Landau levels (see Supplemental Material) [58].
We consider the wave packet dynamics in a harmonic potential that can arise from external electrostatic fields or Coulomb interactions in Wigner crystals [78, 81]
$V(\bm{x})=\frac{\kappa}{2}\bm{x}^{2}.$ (12)
Thus, we have the equation of motion $\dot{q}^{i}=-\kappa x^{i}$ and thereby $\dot{x}^{i}=-\frac{1}{\kappa}\ddot{q}^{i}$. Here $x^{i}=\delta^{ij}x_{j}$ and $\delta^{ij}$ the Kronecker delta. The equation of $\bm{q}$ can be expressed as a forced geodesic equation in the momentum space ($\bm{q}$-space):
$\ddot{q}^{i}+\bar{\Gamma}^{i}_{jk}\dot{q}^{j}\dot{q}^{k}=$ $\bar{G}^{ij}\left(-\frac{\partial E}{\partial q^{j}}+F_{jk}\dot{q}^{k}\right),$ (13)
where $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol that vanishes in this example as $G_{ij}$ is constant. $F_{jk}$ corresponds to the artificial magnetic field in the momentum space from Berry curvature. This equation follows the Lagrangian [58]
$L=\frac{1}{2}\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}-A_{i}\dot{q}^{i}$ (14)
where $A_{j}$ is a vector potential and the band energy is a constant that is set as zero. This Lagrangian has a well-defined canonical structure enabling us to perform canonical quantization. The canonical momentum is $p_{i}=\frac{\partial L}{\partial\dot{q}^{i}}=\bar{G}_{ij}\dot{q}^{j}-A_{i}$ and the Hamiltonian reads
$H=\frac{1}{2M}\delta^{ij}(p_{i}+A_{i})(p_{j}+A_{j})$ (15)
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. By promoting $(p_{i},q^{j})$ to operators with the commutator $\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian in momentum-space representation. We note that here the canonical momentum $\hat{p}_{i}$ corresponds to the real-space coordinate, ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 30, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus [84, 85]. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F\kappa}{1+\kappa G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass, whereas the nonadiabatic metric contributes to a correction from interband mixing. The interband mixing effect is enhanced by increasing the potential strength that increases $\dot{\bm{q}}$ and thus the interband contributions.
Figure 2: Bound states of flat-band particles in real space mirror the Landau-level wave functions on a torus in $\bm{q}$-space.
The confining problem can also be generalized to two-body bound states [30] where the effect of $G_{0}$ can be enhanced as $\kappa$ originates from the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the wave packets are well separated, and they interact through a softened Coulomb interaction in real space. The two-particle Lagrangian is
$L_{\rm T}=L({\bm{x}}_{1})+L({\bm{x}}_{2})+\frac{1}{\epsilon\sqrt{(\bm{x}_{1}-\bm{x}_{2})^{2}+x_{0}^{2}}}$ (16)
where $\bm{x}_{1,2}$ are wave packet centers, $L(\bm{x}_{1,2})$ is the single-particle Lagrangian with the same constant Berry curvature $F$ and nonadiabatic metric $G_{0}$, $\epsilon$ is the effective dielectric constant, the charge is unity, and $x_{0}$ is the cutoff length. Focusing on bound states with $|\bm{x}_{1}-\bm{x}_{2}|\ll x_{0}$, we approximate the Lagrangian as
$L_{\rm T}\simeq L({\bm{x}}_{1})+L({\bm{x}}_{2})+\left[V_{0}-\frac{1}{2}\kappa_{\rm I}({\bm{x}}_{1}-{\bm{x}}_{2})^{2}\right]$ (17)
where $V_{0}=\frac{1}{\epsilon x_{0}}$ and $\kappa_{\rm I}=\frac{V_{0}}{x_{0}^{2}}$.
We focus on the relative dynamics. Applying canonical quantization, the effective Hamiltonian for the relative coordinate is
$\hat{H}_{\rm-}=\frac{1}{2M^{\prime}}(\hat{\bm{p}}_{-}+\hat{\bm{A}}(\hat{\bm{q}}_{-}))^{2}-V_{0}$ (18)
where $\hat{\bm{q}}_{-}=(\hat{\bm{q}}_{1}-\hat{\bm{q}}_{2})/\sqrt{2}$, $\hat{p}_{-,i},\hat{q}_{-}^{j}]=-i\delta_{i}^{j}$, and $M^{\prime}=G_{0}+\frac{1}{2\kappa_{\rm I}}$ is the reduced effective mass. This Hamiltonian again resembles a Landau level problem on a torus (see Fig. 3) with eigenenergies $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M^{\prime}}-V_{0}$.
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. If they are indistinguishable, the two-body wavefunctions need to be antisymmetric for fermions or symmetric for bosons. This leads to another boundary condition in addition to the periodic boundary conditions. Therefore, while the degeneracy of eigen-energy levels obtained by solving equation (18) is the Chern number $|C|$ ($C\neq 0$), only those states satisfying both boundary conditions are allowed solutions.
A natural realization of the two-body problem is the excitons formed by electrons and holes in flat bands [86, 30, 87]. The potential relationship with Cooper pairs is interesting and warrants further investigation [88, 89, 90, 91].
Figure 3: Wave packet dynamics in $\bm{q}$-space, a torus with curved geometry. The distance $ds^{2}$ is defined by the nonadiabatic metric $G_{ij}$. The topological charge stands for the Dirac monopole that generates nonzero gauge fields corresponding to the Berry curvature. $C$ indicates the Chern number. Color encodes the scalar potential from the band dispersion.
II.4 Metric Tensor and Forced Geodesic Equation
The nonadiabatic metric also underscores the importance of the $\bm{q}$-space picture, instead of the real space, where $\bm{q}$ is taken as a generalized coordinate to describe the wave packet dynamics. This enables us to study the emergent gravity and emergent mass in momentum space, which will be discussed in this and following sections, separately.
$G_{ij}(\bm{q})$ in the Lagrangian defines a metric tensor in the tangent vector space of the generalized velocity $\dot{\bm{q}}$ defining a line element $ds^{2}=G_{ij}dq^{i}dq^{j}$. In contrast to the quantum metric, which defines the quantum distance $ds_{Q}$ between two wave functions $|u_{\bm{q}}\rangle$ and $|u_{\bm{q}+d\bm{q}}\rangle$ at $\bm{q}$ and $\bm{q}+d\bm{q}$ via $ds_{Q}^{2}=1-|\langle u_{\bm{q}}|u_{\bm{q}+d\bm{q}}\rangle|^{2}$ [3], the nonadiabatic metric defines the geometric distance between the momentum-space centers of two wave packets. A $\bm{q}$-dependent $G_{ij}$ signifies a curved manifold. Conversely, in cases where $G_{ij}$ is constant across the momentum space, it corresponds to an effective mass in a flat $\bm{q}$-space, simplifying the dynamics to a Euclidean framework.
Expressing the equations of motion in the $\bm{q}$-space clarifies the geometric interpretation. When the inverse of the metric tensor exists, the equation of motion reads
$\ddot{q}^{i}+\Gamma^{i}_{jk}\dot{q}^{j}\dot{q}^{k}=$ $G^{ij}\left(-\frac{\partial E}{\partial q^{j}}+F_{jk}\dot{q}^{k}+\dot{x}_{j}\right)$ (19)
where $\Gamma^{l}_{jk}$ is the Christoffel symbol of the second kind
$\Gamma^{i}_{jk}=\frac{1}{2}G^{il}\left(\frac{\partial G_{lj}}{\partial q^{k}}+\frac{\partial G_{kl}}{\partial q^{j}}-\frac{\partial G_{jk}}{\partial q^{l}}\right),$ (20)
$G^{ji}G_{ik}=\delta^{j}_{k}$ with $\delta^{j}_{k}$ the Kronecker delta. When the right-hand side of equation (19) vanishes, the equation reduces to the geodesic equation for a free particle navigating a curved manifold with metric tensor $G_{ij}$. A nonzero right-hand side indicates a particle subject to forces from three sources (see Fig. 3): a scalar potential from the band dispersion $E$, a gauge field from the Berry curvature $F_{ij}$, and an external force from the wave packet velocity $\dot{x}_{j}$. This equation enables the study of gravity in condensed matter setups [92]. It is noted that equation (19) differs from the recent study on momentum space gravity [93, 94, 95] where $G^{ij}$ and $\ddot{\bm{q}}$ are absent and $\Gamma_{i,jk}$ instead of $\Gamma^{i}_{jk}$ is used.
When $x_{i}$ can be invertibly represented as a function of $\dot{q}^{j}$ by using $\dot{q}^{i}=-\frac{\partial E}{\partial x_{i}}$, we have
$x_{i}=x_{i}(\dot{\bm{q}}),~~~\dot{x}_{i}=\frac{\partial x_{i}}{\partial\dot{q}^{j}}\ddot{q}^{j}=-\Xi_{ij}\ddot{q}^{j},$ (21)
the equations of motion find their $\bm{q}$-space representation:
$\ddot{q}^{i}+\bar{\Gamma}^{i}_{jk}\dot{q}^{j}\dot{q}^{k}=$ $\bar{G}^{ij}\left(-\frac{\partial E}{\partial q^{j}}+F_{jk}\dot{q}^{k}\right),$ (22)
where $\bar{G}_{ij}=G_{ij}+\Xi_{ij}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol. With information of $\dot{\bm{q}}$, $x_{i}$ can be solved using algebraic equations. The metric tensor becomes dynamical with $\dot{\bar{G}}_{ij}=\frac{\partial G_{ij}}{\partial q_{k}}\dot{q}_{k}+\frac{\partial\Xi_{ij}}{\partial\dot{q}_{k}}\ddot{q}_{k}$.
III Discussions
We introduce a geometric framework that extends the study of Bloch electron dynamics in a scalar potential beyond adiabatic Berry-phase effects. We show that the leading-order corrections to the adiabatic dynamics, the nonadiabatic effects, are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor in semiclassical wave packet dynamics, distinguishing them from the conventional quantum metric. This metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature, enabling further generalization. It also modifies the quantum electron dynamics, manifested as a metric in momentum space. A constant nonadiabatic metric influences the flat-band electron dynamics in a harmonic potential by introducing a correction to the effective mass for the generalized coordinate $\bm{q}$. Depending on the origin of the harmonic potential, this mass correction influences the electronic properties in static potentials, Wigner crystal [78, 81], or flat-band excitons [30]. In general, the nonadiabatic metric is not a constant, which endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space. This enables the study of momentum-space gravity.
IV Methods
IV.1 Integrate out $c_{n}$
Given the Lagrangian
$L$ $=\sum_{n}c_{n}^{*}i\dot{c}_{n}-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}+\Lambda_{[00]}+\sum_{n\neq 0}-c_{n}^{*}c_{n}\left[\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]})\right]+(c_{0}^{*}c_{n}\Lambda_{[0n]}+c.c.)$
where $c_{0}$ is the reference-band amplitude and $c_{n}$ are dynamical variables similar to $\bm{x}_{c}$ and $\bm{q}_{c}$. The normalization implies $c_{0}=1+\mathcal{O}(c_{n\neq 0}^{2})$ and the phase of $c_{0}$ contributes a total derivative and can be gauged away. For simplicity, we define
$\Delta_{n}\equiv\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]}),\qquad J_{n}\equiv c_{0}\Lambda_{[n0]},~n\neq 0.$ (23)
The action $S=\int dtL$ is quadratic in the excited-band amplitudes $c_{n\neq 0}$. Varying S with respect to $c_{n}^{*}$ gives the saddle-point equation
$(i\partial_{t}-\Delta_{n})c_{n}=-J_{n}\qquad n\neq 0.$ (24)
[…]
$S_{\rm eff}[\bm{q}_{c},\bm{x}_{c}]$ $=\int dtL_{0}(\bm{q}_{c},\bm{x}_{c})$
$-\sum_{n\neq 0}\int dtdt^{\prime}J_{n}^{*}(t)G_{n}(t,t^{\prime})J_{n}(t^{\prime}),$ (25)
where $L_{0}(\bm{q}_{c},\bm{x}_{c})\equiv-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c,0}+\Lambda_{[00]}$.
When parameters $\bm{q}_{c}$ vary slowly, we expand the inverse operator:
$G_{n}$ $=(i\partial_{t}-\Delta_{n})^{-1}\delta(t-t^{\prime})$
$=\left[-\frac{1}{\Delta_{n}}+\frac{i}{\Delta_{n}^{2}}\partial_{t}+\mathcal{O}(\partial_{t}^{2})\right]\delta(t-t^{\prime}),$ (26)
and keep terms up to second order in time derivatives. Using $J_{n}=c_{0}\Lambda^{n0}$ and integrating by parts, equation (25) gives a local effective Lagrangian
$\mathcal{L}_{\rm eff}=L_{0}+\sum_{n\neq 0}\frac{J_{n}^{*}J_{n}}{\Delta_{n}}+\mathcal{O}(\partial_{t}).$ (27)
By substituting $J_{n}=c_{0}\dot{\bm{q}}_{c}\cdot\bm{A}_{[n0]}$ into the above equation and neglecting the Berry phase corrections to the band gap, i.e., $|\Lambda_{[nn]}-\Lambda_{[00]}|\ll|\omega_{[n0]}|$ and $\Delta_{n}\simeq\omega_{[n0]}$, we obtain the effective Lagrangian as shown in equation (6).
V Acknowledgments
Y.R. thanks Qian Niu, Yang Gao, Di Xiao, Wenqin Chen and Jie Wang for useful discussions.